OpenStudy (anonymous):

Solve: (x-6/x)-1 = x+4/x

OpenStudy (anonymous):

\[\frac{ x-6 }{ x }-1=\frac{ x+4 }{ x }\]

OpenStudy (harsimran_hs4):

keep a check x cannot be zero multiply both sides by x and then obtain quadratic eq and solve

OpenStudy (leksi):

x=-10

OpenStudy (anonymous):

\[x \times \frac{ x-6 }{ x }-1=\frac{ x+4 }{x }\times x\]

OpenStudy (luigi0210):

don't forget to multiply the 1 too

OpenStudy (anonymous):

oh okay so x-6-1x=x+4 is the quad. equation? And thank you Leksi but I want to know how to figure it out...

OpenStudy (luigi0210):

Yup, now move everything to one side

OpenStudy (anonymous):